OpenStudy (anonymous):

in order to determine her pulse rate, rate jimes nurse counted 18 beats in her pulse in 15 seconds at this rate how many beats would she havein 60 seconds

OpenStudy (anonymous):

WELCOME TO OPEN STUDY!! To find the answer you could be 60/15=4 then multiply 18(4). Welcome again!!

OpenStudy (mathstudent55):

You can use a proportion: \(\dfrac{18~beats}{15~sec} = \dfrac{b~beats}{60~sec} \) Solve for b.

OpenStudy (anonymous):

Id trust these people's answer more then my answer :D

OpenStudy (mathmale):

\[\frac{ 18~beats }{ 15~seconds }\]Use the conversion factor Now multiply '60 seconds' by this conversion factor. Notice that the result will have the unit of measurement "beats."